OpenStudy (anonymous):

What is the vertex of the function f(x)=2x^2+4x-16?

OpenStudy (anonymous):

(-2,4) ?

OpenStudy (jhannybean):

standard form y = a(x-h)^2+k y= 2x^2 + 4x - 16 move the 16 to the other side y + 16 = 2x^2 + 4x factor out a 2 from the right side. y + 16 = 2(x^2 + 2x) complete the square y + 16 + 2( 1 ) = 2( x^2 +2x + 1) simplify. y + 18 = 2(x + 1)^2 y = 2(x + 1)^2 - 18 put it back in standard form. y = 2( x - (-1))^2 + (-18)

OpenStudy (anonymous):

Refer to the attached Mathematica calculation.

OpenStudy (jhannybean):

Am i wrong? :\

OpenStudy (jhannybean):

Oh i see.

OpenStudy (jhannybean):

WELL!! we got the same thing, lol. (-1,-18) for the vertex. Thanks @robtobey :)